OpenStudy (anonymous):

1+cos2thita=2cossqare thita solve plzzz

OpenStudy (aravindg):

\[\cos 2 \theta= \cos^2 \theta-\sin^2 \theta\]

OpenStudy (aravindg):

\[1=\sin^2 \theta +\cos ^2 \theta \]

OpenStudy (anonymous):

then

OpenStudy (aravindg):

now cant u do it ? :)

OpenStudy (aravindg):

substitute in LHS to get RHS hence proved !